# Magnon thermal Hall effect in collinear ferrimagnets

### v3

@@ Abstract
therefore, a combination $\bar{w}_{xy;{\bf k}}T_{\bf k}$ of the Berry curvature has a part $\propto t\sin(\zeta+2\theta)\cos^{2}\left(\frac{\sqrt{3}k_{x}}{2}\right)\sin^{2}\left(\frac{k_{y}}{2}\right)$ which doesn’t vanish upon angle integration. We can conclude that THE will be non-zero if there is no symmetry between the two magnetic sublattices. Certain anisotropic HEI and DMI appear in the Hamiltonian of magnons as a result of broken symmetry between the sublattices. For example, in our model, amplitude of THE is defined by such anisotropic HEI and DMI within the sublattices or anisotropic HEI within the sublattices and DMI between the sublattices. We checked that there is no THE when $t=\zeta=0$ and $\theta\neq 0$, meaning that DMI between the sublattices can be gauged away from the dispersion of magnons in this case. Namely, although such a DMI seemingly may make the Berry curvature Eq. (23) not to vanish upon angle integration, a simple shift of the momentum $\frac{k_{y}}{2}-\theta\rightarrow\frac{k_{y}}{2}$ removes $\theta$ from the Berry curvature making THE to vanish. However, such a shift doesn’t fully gauge away the DMI $\theta$ when $\cos\left(\frac{k_{y}}{2}+\zeta\right)$ is present in Eq. (24) thus making THE to be non-zero. We plot $\sigma^{\mathrm{THE}}$ defined as $j^{\mathrm{TH}}_{\alpha}=\sigma^{\mathrm{THE}}\nabla_{\beta}\chi({\bf r})$ in the left of Fig. (3) for $\theta=0.2$, $t=0.4$ and $\chi=0$ as a function of temperature. We note that HEI within the sublattices can be isotropic (the green atom doesn’t make $t=0$ for the interaction across it) in our model for non-zero THE. However, it is of the two orders of magnitudes less than that shown in the left of Fig. (3).
@@
To conclude, we have studied a theoretical model of an insulating collinear antiferromagnet which shows magnon thermal Hall effect. We showed that for the magnon thermal Hall effect to be non-zero, the symmetry between the sublattices of an antiferromagnet must be broken. The antiferromagnet in this case is a ferrimagnet. This is in contrast to Dzyaloshinskii’s weak ferromagnets Dzyaloshinskii1958 where magnon thermal Hall is also expected. We have corroborated our results by symmetry argument consistent with Dzyaloshinskii’s classification scheme of antiferromagnets Turov1965 . In addition, we found magnon spin-splitting at the $\bm{\Gamma}$} point due to the Dzyaloshinskii-Moriya to be one of the ingredients for non-zero magnon thermal Hall effect. We suggested a scenario when external electric field applied in the plane of the system can change magnon thermal Hall effect by altering symmetry of the lattice.

### v4

@@ V Ferrimagnet
therefore, a combination $\bar{w}_{xy;{\bf k}}T_{\bf k}$ of the Berry curvature has a part $\propto t\sin(\zeta+2\theta)\cos^{2}\left(\frac{\sqrt{3}k_{x}}{2}\right)\sin^{2}\left(\frac{k_{y}}{2}\right)$ which doesn’t vanish upon angle integration. We can conclude that THE will be non-zero if there is no symmetry between the two magnetic sublattices. Certain anisotropic HEI and DMI appear in the Hamiltonian of magnons as a result of broken symmetry between the sublattices. For example, in our model, amplitude of THE is defined by such anisotropic HEI and DMI within the sublattices or anisotropic HEI within the sublattices and DMI between the sublattices. We checked that there is no THE when $t=\zeta=0$ and $\theta\neq 0$, meaning that DMI between the sublattices can be gauged away from the dispersion of magnons in this case. Namely, although such a DMI seemingly may make the Berry curvature Eq. (23) not to vanish upon angle integration, a simple shift of the momentum $\frac{k_{y}}{2}-\theta\rightarrow\frac{k_{y}}{2}$ removes $\theta$ from the Berry curvature making THE to vanish. However, such a shift doesn’t fully gauge away the DMI $\theta$ when $\cos\left(\frac{k_{y}}{2}+\zeta\right)$ is present in Eq. (24) thus making THE to be non-zero. We plot $\sigma^{\mathrm{THE}}$ defined as $j^{\mathrm{TH}}_{\alpha}=\sigma^{\mathrm{THE}}\nabla_{\beta}\chi({\bf r})$ in the left of Fig. (4) for $\theta=0.2$, $t=0.4$ and $\chi=0$ as a function of temperature. We have checked that finite $\chi$ also results in the THE of magnons.
@@
VI Collinear weak ferromagnet
Having understood genuine antiferromagnet and ferrimagnetic systems, we now turn our attention to the weak ferromagnetBorovikRomanov1957 ; Dzyaloshinskii1958 case. Weak ferromagnet is a Néel ordered antiferromagnet in which the symmetry allows for a finite magnetic moment. Finite magnetic moment might be due to the canting of the Néel order or be carried by magnons (or fermions in case the system is metallic). In a weak ferromagnet Dzyaloshinskii1958 the two magnetic sublattices are connected to each other by some symmetry operation, while in the ferrimagnet the two magnetic sublattices aren’t. We consider the case when the Néel order is collinear, i.e. there is no canting of the order, and the magnetic moment is carried by the magnons. We stress that canting of the Néel order isn’t necessary for an antiferromagnet to be a weak ferromagnet. This is because it is rather the symmetry that distinguishes weak ferromagnets from other collinear antiferromagnets Dzyaloshinskii1958 ; Turov1965 .
Figure 5: The green atom is lifted from the xyx-y plane making the system a weak ferromagnet with the $M_{z}L_{x}$ Dzyaloshinskii’s invariant. Purple arrows are the directions of the DMI the green atom creates only due to the lifting from the xyx-y plane, ${\bf d}_{1}=\frac{1}{2}(-\sqrt{3},-1)$, ${\bf d}_{2}=\frac{1}{2}(\sqrt{3},-1)$, ${\bf d}_{3}=(0,-1)$. In the right figure only the $x$- component of the DMI, that is relevant for the Néel order in $x$- direction, is shown.
Figure 6: THE of magnons for the weak ferromagnet described by Eq. (25) plotted for $\theta=0.2$, $t=0.4$, $\zeta=0$.
A weak ferromagnetBorovikRomanov1957 ; Dzyaloshinskii1958 can be obtained from the genuine antiferromagnet, shown in the left of Fig. (1), by lifting the green atom from the xyx-y plane. Lifting of the green atom allows for the $M_{z}L_{x}$ Dzyaloshinskii’s invariant in the system. We then set the Néel order in $x$- direction and take into account only such DMI that are activated by a comination of the order in $x$- direction and lifted green atoms. Such DMI is shown in Fig. (5) by ${\bf d}_{1}=\frac{1}{2}(-\sqrt{3},-1)$, ${\bf d}_{2}=\frac{1}{2}(\sqrt{3},-1)$, ${\bf d}_{3}=(0,-1)$. See Ref. KovalevZyuzinLiPRB2017 for an example of how DMI is defined with these vectors. When the Néel order is in $x$- direction, only the $x$ component of ${\bf d}_{1/2/3}$ is relevant, and we can reduce the DMI only to these components. The plus and minus signs in the right figure in Fig. (4) is the reduced DMI. The exchange Hamiltonian between the neighboring spins and the DMI due to the lifting of the green atom from the xyx-y plane is
$H_{\mathrm{weak}}=\sum_{\langle ij\rangle}$ $\left[JS_{i}^{x}S_{j}^{x}+J\cos(\theta_{ij})\left(S_{i}^{y}S_{j}^{y}+S_{i}^{z}S_{j}^{z}\right)\right]$
$+\sum_{\langle ij\rangle}J\sin(\theta_{ij})\left(S_{i}^{y}S_{j}^{z}-S_{i}^{z}S_{j}^{y}\right).$ (25)
We thus have magnons described by Eq. (7), in which $\gamma_{\bf k}$ is defined in Eq. (9) and $t^{\mathrm{A}/\mathrm{B}}_{\bf k}$ are defined in Eqs. (10) and (11). Therefore, magnons describing weak ferromagnet are a mixture of those describing genuine antiferromagnet and ferrimagnet. We stress that in the description of magnons the DMI in the weak ferromagnet is identical to the ferrimagnetic case. THE of magnons for the weak ferromagnet Eq. (25) is plotted in Fig. (6). There are more weak ferromagnets in Nature than genuine antiferromagnets Turov1965 . This maybe understood from the examples used in this paper. Genuine antiferromagnet given by the model A shown in the left of Fig. (1) is unstable towards formation of the weak ferromagnet because the in-plane position of the green atom is rather a fine tuned situation. Indeed, even a slight lifting of the position of the green atom from the plane of the crystal will turn genuine antiferromagnet into a weak ferromagnet. Realistic weak ferromagnets are presented by RuO2, CrSb, CoF2, NiF2, $\alpha$-Fe2O3, MnTe, LuFeO3, and many more. If one wants to undertand whether a given antiferromagnet belongs to the symmetry class of a weak ferromagnet, one has to look for the symmetry group of the system at hand and consult the table 3 in Turov1965 . Presence or absence of the magnetic moment in these materials depends on the orientation of the Néel order in accord with the Dzyaloshinskii’s invariant. For example, in the studied here system shown in Fig. (5) only the $L_{x}$ can generate finite magnetic moment $M_{z}$. Clearly, THE of magnons will be absent if the magnetic moment isn’t allowed for certain directions of the Néel order.
VII Conclusions
@@
To conclude, we have studied theoretical models, depicted in Fig. (1) and Fig. (5), of insulating compensated (total spin on the two sublattices is zero) collinear antiferromagnets that show magnon thermal Hall effect. We showed that for the magnon thermal Hall effect to be non-zero, the symmetry between the sublattices of an antiferromagnet must be broken. The antiferromagnet in this case is a ferrimagnet. We then showed that finite magnon thermal Hall effect can occur also in the case when the two magnetic sublattices are connected by some symmetry operation. The antiferromagnet in this case is a collinear weak ferromagnet Dzyaloshinskii1958 ; Turov1965 . We have corroborated our results by symmetry argument consistent with Dzyaloshinskii’s classification scheme of antiferromagnets Turov1965 . In particular we have deduced Dzyalshinskii’s invariant for every studied systems and showed that finite magnon thermal Hall effect in ferrimagnet and collinear weak ferromagnet is consistent with the invariants. In addition, we found magnon spin-splitting at the $\bm{\Gamma}$} point due to the Dzyaloshinskii-Moriya to be one of the ingredients for non-zero magnon thermal Hall effect. We suggested a scenario when external electric field applied in the plane of the system can change magnon thermal Hall effect by altering symmetry of the lattice.